OpenStudy (anonymous):

\[\Huge\int\dfrac{dx}{4+2^x}\]

OpenStudy (goformit100):

How you posted this Question ?

OpenStudy (anonymous):

What do you mean? I just type in question and open it?

OpenStudy (goformit100):

But how did you used Mathematics Equation way in posting question

OpenStudy (anonymous):

sorry, still dont understand what you are asking? I'm using Latex, if this is what you want to know

OpenStudy (loser66):

Let u = 2^x

OpenStudy (kainui):

Why are you using latex and not like cotton or polyester?

OpenStudy (goformit100):

How did you got latex before posting your question

OpenStudy (anonymous):

I just use `\[ anything here \]`...

OpenStudy (anonymous):

ok so \(u = 2^x\) and \(du = 2^x\ln 2 dx\), right? but where to sub in for dx?

OpenStudy (kainui):

I'd suggest using u=4+2^x instead.

OpenStudy (anonymous):

u = 4+2^x

OpenStudy (anonymous):

ye

OpenStudy (goformit100):

But this dosent come when I post my questions

OpenStudy (anonymous):

but there is no 2^xln2 dx to sub

OpenStudy (kainui):

Well, actually I don't think it matters to be honest, @Loser66 's substitution is just as good.

OpenStudy (anonymous):

\[\int\dfrac{dx}{u}\]??

OpenStudy (loser66):

u =2^x --> du = 2^x ln2 dx--> dx = \(\dfrac{du}{2^xln2}=\dfrac {du}{uln2}\)

OpenStudy (loser66):

ok then, right?

OpenStudy (anonymous):

wow so thats how you sub, eh? I always though i have to sub both u and du. cool

OpenStudy (kainui):

Yeah, pretty spiffy.

OpenStudy (anonymous):

So \[\dfrac{1}{\ln 2}\int\dfrac{du}{u} = \dfrac{\ln|4+2^x|}{\ln2}+C\]

OpenStudy (loser66):

???

OpenStudy (loser66):